# SLOPE and Designing Robust Studies for Generalization

### v1

@@ Abstract
A popular task in generalization is to learn about a new, target population based on data from an existing, source population. This task relies on conditional exchangeability, which asserts that differences between the source and target populations are fully captured by observable characteristics of the two populations. Unfortunately, this assumption is often untenable in practice due to unobservable differences between the source and target populations. Worse, the assumption cannot be verified with data, warranting the need for robust data collection processes and study designs that are inherently less sensitive to violation of the assumption. In this paper, we propose SLOPE (Sensitivity of LOcal Perturbations from Exchangeability), a simple, intuitive, and novel measure that quantifies the sensitivity to local violation of conditional exchangeability. SLOPE combines ideas from sensitivity analysis in causal inference and derivative-based measure of robustness from Hampel, (1974). Among other properties, SLOPE can help investigators to choose (a) a robust source or target population or (b) a robust estimand. Also, we show an analytic relationship between SLOPE and influence functions (IFs), which investigators can use to derive SLOPE given an IF. We conclude with a re-analysis of a multi-national randomized experiment and illustrate the role of SLOPE in informing robust study designs for generalization.
@@ 1 Introduction
1.1 Motivation and Overview
There has been a growing interest in generalizing information from a source population to a new, target population under the assumption of conditional exchangeability, which posits that distributions in the two populations are identical after conditioning on a set of populations’ shared characteristics. Formally, suppose each datum is represented as random variables $(O,X)$ where $X$ denotes the shared characteristics between the two populations. The goal is to learn the target distribution $Q_{O,X}$ given (i) the “full” data $(O,X)$ from the source distribution $P_{O,X}$ and (ii) the “partial” data $X$ about the shared characteristics in the target distribution $Q_{X}$. Conditional exchangeability states
@@ 1.1 Motivation and Overview
Under (1), we can learn the target distribution $Q_{O,X}$ via a heuristic yet simple equality
@@
The first equality is from the definition of conditional probability and the second equality is from (1). Also, when we are interested in a feature/functional of the target distribution $Q_{O,X}$, denoted as $\psi(Q_{O,X})$ and referred to as the target estimand (e.g., mean or median of $O$ in the target population), we can identify the estimand under (1) by $\psi(Q_{O,X})=\psi(P_{O\mid X}\times Q_{X})$. For concrete examples, see Section 2.1.
Unfortunately, recent works (Allcott,, 2015; Jin et al.,, 2024) have argued that conditional exchangeability is likely violated in practice, potentially threatening the validity of several empirical works on generalization and transfer learning. Worse, conditional exchangeability is inherently untestable because we never observe $O$ from the target distribution $Q$ (Dahabreh et al.,, 2023; Huang,, 2024); as mentioned above, we only know $P_{O,X}$ and $Q_{X}$. Taken together, before analyzing data for generalization, it’s essential to have robust data collection processes and study designs that are inherently less sensitive to violation of the assumption.
To this end, this paper proposes a simple, interpretable, and novel tool that helps gauge which study designs are robust to violations of conditional exchangeability and we present a high-level summary of the tool and its important caveats. Let $Q_{O\mid X}^{\gamma}$ be the distribution of $Q_{O\mid X}$ when conditional exchangeability is violated by a degree quantified by a sensitivity parameter $\gamma\in\mathbb{R}$ and $\gamma=0$ implies conditional exchangeability holds (i.e., $Q_{O\mid X}^{0}=P_{O\mid X}$). Then, we propose a metric called SLOPE, which stands for Sensitivity of LOcal Perturbations from Exchangeability:
@@
Figure 1 provides a visual illustration of SLOPE. As the name hints, SLOPE is the slope of the tangent of the target estimand $\psi(Q_{O,X}^{\gamma})$ as a function of $\gamma$ at $\gamma=0$. In other words, SLOPE measures the change in the target estimand under $Q_{O,X}^{\gamma}$ (i.e., when conditional exchangeability doesn’t hold) and $Q_{O,X}^{0}$ (i.e., when conditional exchangeability holds) when there is a near-violation of conditional exchangeability. Generally speaking, a higher magnitude of SLOPE suggests that the estimand changes more dramatically to local violations of conditional exchangeability and a lower magnitude of SLOPE suggests otherwise. More broadly, SLOPE aligns with principles in robust statistics, where “robustness signifies insensitivity to small deviations from the assumptions” (Huber,, 1981, p. 1), where we added the emphasis on “small deviations”.
@@
SLOPE depends on two quantities: (a) the target estimand of interest (i.e., $\psi$) and (b) the target distribution under conditional exchangeability (i.e., $Q_{O,X}^{0}=P_{O\mid X}\times Q_{X}$). Importantly, SLOPE does not depend on choices of any estimation procedure of $\psi$. To put it differently, SLOPE is an intrinsic, design-level characteristic about (a) the estimand/functional (i.e., $\psi$) and (b) the distributions that are involved in generalization (i.e., $P_{O\mid X}$ and $Q_{X}$) when there is a violation of conditional exchangeability.
@@
The simplicity of SLOPE comes at some costs, which we briefly summarize here and discuss in detail in Section 3.1. First, as mentioned above, SLOPE is a local measure and for small deviations of $\gamma$ from 0, SLOPE is an accurate reflection about the change in $\psi$. For larger deviations (i.e., $\gamma$ is far from 0), SLOPE may still provide valuable intuition, subject to the usual limitations of linear approximations based of the tangent line. We also remark that some well-known robustness measures are local, including Hampel, (1974)’s influence function, which has taken a central role in robust statistics and estimation (e.g., for causal parameters). Second, the unit of SLOPE inherits the unit of the target estimand. While this fact is rather obvious from SLOPE’s definition since SLOPE measures the change in the target estimand, which generally has scientifically meaningful units, when interpreting the magnitude of SLOPE, we remind investigators to exercise the same level of caution as when they interpret the units of slopes in other areas of statistics (e.g., regression models). Third, SLOPE depends on the parametrization of $Q_{O\mid X}^{\gamma}$ or equivalently, the sensitivity model for conditional exchangeability. While our parametrization of $Q_{O\mid X}^{\gamma}$ has some technical and non-technical benefits, for instance (a) analytic derivation of SLOPE only requires basic tools from calculus, (b) the uniqueness of the tangent curve, especially compared to “bound-based” sensitivity models (see Remark 3), and (c) interpretability (see Section 3), it is not the only parametrization of $Q_{O\mid X}^{\gamma}$; see Rosenbaum, (1987), Tan, (2006), Ding and VanderWeele, (2016), Oster, (2019), and Cinelli and Hazlett, (2020) for other models of sensitivity analysis. Irrespective of the choice of the sensitivity model, we believe the high-level idea of SLOPE as a simple, derivative-based summary of a sensitivity analysis can provide new and important insights about designing robust studies for generalization.
@@
Figure 1: Left: the data structure of our setup. Right: a visual illustration of SLOPE. The $y$-axis plots the target estimand $\psi(\cdot)$ and the x-axis plots $\gamma$ which represents the violation of conditional exchangeability; $\gamma=0$ indicates that conditional exchangeability holds.
@@ 1.2 Prior Works
As mentioned in our summary above, SLOPE and Hampel, (1974)’s influence function (IF) are related in that both SLOPE and IF measure local sensitivity (or robustness) and use derivatives to quantify sensitivity (robustness). But, there is an important difference between the two measures. Specifically, Hampel, (1974)’s IF measures a local change in an estimator due to contamination of a single data point whereas SLOPE measures a local change in an estimand due to violation of conditional exchangeability. Despite these differences, SLOPE and IF are analytically related. Section 4.1 formalizes this insightful relationship between the two local measures of robustness.
@@
Our work fits into the broader literature on sensitivity analysis of conditional exchangeability for generalizability and transportability (Nguyen et al.,, 2017; Nie et al.,, 2021; Colnet et al.,, 2021; Dahabreh et al.,, 2022, 2023; Duong et al.,, 2023; Ek and Zachariah,, 2023; Huang,, 2024; Jin et al.,, 2024). Notably, our paper uses the exponentially tilting sensitivity model in Dahabreh et al., (2022, 2023), which was originally proposed by Robins et al., (2000) and used by others (e.g., Rotnitzky et al., (2001); Troxel et al., (2004); Birmingham et al., (2003); Franks et al., (2020); Nabi et al., (2024); Zhou and Yao, (2023)). What differentiates our paper from existing works on sensitivity analysis in this area is that existing works studied sensitivity of estimators of $\psi$. In contrast, our work studies sensitivity of study designs; as mentioned in our summary above, SLOPE is agnostic to how $\psi$ is estimated and is a population-level characteristic.
@@
Outside of generalizability and transportability, there are several works on summary measures about the violation of key assumptions in their respective fields and we highlight some works. Gupta and Rothenhäusler, (2023) proposed the directional s-value which quantifies the minimum amount of covariate shift that alters the sign of the target parameter (usually the mean). Rosenbaum, (2004) proposed design sensitivity $\widetilde{\Gamma}\in(1,\infty)$, which is a scalar, odds-ratio based summary of a test statistic’s power under unmeasured confounding in an observational study. Andrews et al., (2017) considered local mis-specification of generalized method of moments (GMM) and their matrix-based measure, denoted as $\Lambda\in\mathbb{R}^{p\times p}$, quantified changes of the estimators for GMM parameters. In missing data, Troxel et al., (2004) proposed a measure, which quantifies the change in an estimator to a local violation of ignorable missingness. Troxel et al., (2004)’s measure is, loosely speaking, the “estimator” analog of SLOPE in missing data. Oster, (2019), Ding and VanderWeele, (2016), Cinelli and Hazlett, (2020), and Zhao, 2019b proposed scalar measures that summarize the impact of unmeasured confounding in observational studies. Specifically, each proposed metrics that quantify the “minimum unmeasured confounding bias” necessary to alter the original study conclusion under conditional exchangeability (e.g., the sensitivity bound does not contain the null effect). All these works, except for Gupta and Rothenhäusler, (2023) and Rosenbaum, (2004) to some extent, measure robustness of estimators or tests rather than that of estimands or study designs.
Finally, another related line of work assesses generalizability by comparing shared characteristics (i.e., $X$) of the two populations (Stuart et al.,, 2011; Tipton,, 2014; Chen et al., 2023b, ). These works are developed under the assumption of conditional exchangeability, whereas we assess robustness when the assumption is subject to violation. Also, while these works focus on generalizing the average treatment effect, our framework is applicable to a wide range of functionals/estimands that include the average treatment effect as a special case.
1.3 Organization and Notation
@@
The rest of the manuscript is organized as follows. Section 2 introduces the setup and the sensitivity model. Section 3 formally introduces SLOPE, its properties, examples, and illustrations on using SLOPE for robust study designs. Section 4 discusses further insights and results, including the relationship between SLOPE and IF, estimation of SLOPE, and other sensitivity models. Section 5 demonstrates the application of SLOPE in robust designs through a re-analysis of a multi-site experiment by Banerjee et al., (2015). Section 6 concludes with a discussion on future directions and practical considerations. Technical proofs and auxiliary results are deferred to the Supplement.
@@
We define the notations that we use throughout the paper. For a probability measure (i.e., a population) $P$ and random vectors $X_{1}$, $X_{2}$, we let $P_{X_{j}}$ be the marginal distribution of $X_{j}$ for $j=1,2$. Given $P_{X_{j}}$, we let $\textnormal{E}_{P_{X_{j}}}(\cdot)$ and $F_{P_{X_{j}}}$ be the expectation and cumulative distribution function (c.d.f.), respectively, under $P_{X_{j}}$. We let $P_{X_{1}\mid X_{2}}(\cdot\mid\cdot)$ be the conditional distribution of $X_{1}$ given $X_{2}$. Similarly, we let $\textnormal{E}_{X_{1}\mid X_{2}}(\cdot\mid\cdot)$, $F_{X_{1}\mid X_{2}}(\cdot\mid\cdot)$, $f_{X_{1}\mid X_{2}}(\cdot\mid\cdot)$ be the conditional distribution, conditional c.d.f., and conditional probability density function, respectively. We also let $\textnormal{E}_{X_{1}\mid X_{2}}(\cdot\mid x_{2})$ be the conditional expectation given a specific $X_{2}=x_{2}$. Throughout the paper, we assume sufficient regularity conditions for conditional distributions and conditional densities to exist; see Shao, (2008, Chapter 1) for the regularity conditions. Finally, for a probability distribution $Q$ defined on the same measurable space of $(X_{1},X_{2})$, we use $P_{X_{1}\mid X_{2}}\times Q_{X_{2}}$ to denote the joint distribution where $P_{X_{1}\mid X_{2}}\times Q_{X_{2}}(A\times B)=\int_{B}P_{X_{1}\mid X_{2}}(A\mid x_{2})dQ_{X_{2}}(x_{2})$ for $A\in\mathcal{B}$ and $B\in\mathcal{B}^{d}$ and $\mathcal{B}$ are Borel $\sigma$-fields with respect to the reals.
@@ 2 Setup
2.1 Goal and Key Assumptions in Generalization
Let $P_{O,X}$ and $Q_{O,X}$ be joint distributions of random variables $(O,X)$ from a source population $P$ and a target population $Q$, respectively, where $O$ is a scalar and $X$ is a vector. Suppose $\psi(\cdot)$ is a scalar-valued functional of distributions that extracts a quantity of scientific interest from the distribution. The goal is to learn the functional (i.e., $\psi$) of the target distribution $Q_{O,X}$, which we denote as $\psi(Q_{O,X})$ and refer to as the target estimand, from the (a) “full” data $(O,X)$ from the source distribution $P_{O,X}$ and (b) “partial” data $X$ from the target distribution $Q_{X}$. See Figure 1 for a visual illustration of the data setup.
We make two remarks on our setup. First, while our exposition considers a scalar functional $\psi$ (e.g., means, medians, average potential outcomes), most of our results extend to a vector valued $\psi$; see Section J.1 of the Supplement for a discussion. Moreover, with a user specified vector norm, the practical interpretation of SLOPE based on its magnitude will also extend; see Section 3.1. Second, almost all results below are agnostic to how the data were sampled within each population, for instance by simple random sampling, i.i.d. sampling, or even adaptive sampling. Specifically, our results remain at the population level until Section 4.2, where we propose an estimation procedure for SLOPE.
Under the setup, the two most popular assumptions to identify the target estimand $\psi(Q_{O,X})$ (e.g., Cole and Stuart, (2010); Tipton, (2014); Kern et al., (2016); Dahabreh et al., (2019); Huang et al., (2023); Zeng et al., (2023); Degtiar and Rose, (2023)) are as follows.

### v2

@@ Abstract
A common task in generalization is to learn about a new target population using data from another source population. This task relies on conditional exchangeability, which assumes that differences between the source and target populations are fully captured by observable variables. However, this assumption is often untenable in practice due to remaining, unobservable differences, and it cannot be verified with data. These limitations warrant the development of robust study designs that are inherently less sensitive to violations of the assumption. We propose SLOPE (Sensitivity of LOcal Perturbations from Exchangeability), a simple and novel measure that quantifies sensitivity to local violations of conditional exchangeability. SLOPE combines ideas from sensitivity analysis in causal inference and derivative-based robustness measure from Hampel’s influence function. To the best of our knowledge, SLOPE is the first metric to quantify the robustness of study designs with respect to violations of conditional exchangeability. Specifically, SLOPE measures the sensitivity of two design-level characteristics: (a) the functional of interest (e.g., the mean or the median) and (b) the study distributions. We demonstrate how SLOPE can guide robust study designs through a re-analysis of a multinational randomized experiment.
@@ 1 Introduction
1.1 Background and Overview
There has been a growing interest in generalizing or transporting information from an existing source population to a new, target population under the assumption of conditional exchangeability and the setup can be formalized as follows. Suppose each datum is represented as random variables $(O,X)$ and the goal is to learn the target distribution $Q_{O,X}$ given (i) the “full” data $(O,X)$ from the source distribution $P_{O,X}$ and (ii) the “partial” data $X$ from the target distribution $Q_{X}$. Conditional exchangeability states that conditional on $X$, the distribution of $O$ between the target and the source population is identical:
@@ 1.1 Background and Overview
Equation (1) enables learning about the target distribution $Q_{O,X}$ based only on (i) $P_{O,X}$ and (ii) $Q_{X}$ and this phenomenon can be illustrated with a heuristic, yet simple equality:
@@
The first equality is from the definition of conditional probability and the second equality is from (1). By the same heuristic argument, we can learn a low-dimensional feature of the target distribution $Q_{O,X}$, denoted as $\psi(Q_{O,X})$ and referred to as the target estimand or target functional, by $\psi(Q_{O,X})=\psi(P_{O\mid X}\times Q_{X})$. Some popular target estimands include the mean of $O$ or the average treatment effect in the target population; see Section 2.1 for details and more examples.
Unfortunately, recent works (Allcott,, 2015; Jin et al.,, 2024) argued that conditional exchangeability is likely violated in practice due to unobservable differences between the source and the target population. Worse, conditional exchangeability is inherently untestable because the variable $O$ is unavailable from the target distribution; under the setup above, we only have access to samples from (i) the joint distribution in the source, $P_{O,X}$, and (ii) the marginal distribution in the target, $Q_{X}$ (Dahabreh et al.,, 2023; Zeng et al.,, 2023; Huang,, 2024). Taken together, these challenges underscore the need to have data collection processes and more broadly, study designs that are inherently less sensitive to violation of conditional exchangeability before analyzing data for generalization.
To this end, the main contribution of the paper is to propose a simple and novel tool that helps gauge which study designs are robust to violations of conditional exchangeability and we present a high-level summary of the tool. Let $Q_{O\mid X}^{\gamma}$ be the distribution of $Q_{O\mid X}$ when conditional exchangeability is violated by a degree quantified by a sensitivity parameter $\gamma\in\mathbb{R}$ and let $\gamma=0$ be the case where conditional exchangeability holds (i.e., $Q_{O\mid X}^{0}=P_{O\mid X}$). Then, we propose a metric called SLOPE, which stands for Sensitivity of LOcal Perturbations from Exchangeability:
@@
As its name and definition imply, SLOPE is the slope of the target estimand $\psi(Q_{O,X}^{\gamma})$ at $\gamma=0$ (see Figure 1 for a visual illustration). SLOPE measures how dramatically the target estimand changes when moving from a setting with no violation of conditional exchangeability (i.e., $\gamma=0$) to a setting with a near-violation (i.e., $\gamma\to 0$). Generally, a higher magnitude of SLOPE suggests that the target estimand is more sensitive/less robust to local violations, while a lower magnitude suggests the estimand is less sensitive/more robust; see Section 3.1 for further discussions on interpreting SLOPE.
@@
SLOPE depends on two quantities: (a) the target estimand (i.e., $\psi$) and (b) the target distribution under conditional exchangeability (i.e., $Q_{O,X}^{0}=P_{O\mid X}\times Q_{X}$). Importantly, SLOPE does not depend on the estimation procedure of $\psi$. To put it differently, SLOPE is an intrinsic, design-level characteristic about (a) the target estimand (i.e., $\psi$) and (b) the source or target distributions in the setup (i.e., $P_{O\mid X}$ and $Q_{X}$) when there is a local violation of conditional exchangeability.
@@
We briefly answer three common and important questions about SLOPE; see Sections 3 and 4 for detailed discussions. First, SLOPE is a local measure and for small deviations of $\gamma$ from 0, SLOPE provides an accurate reflection about the change in $\psi$. For larger deviations of $\gamma$ from 0, SLOPE may still provide valuable intuition, subject to the usual limitations of linear approximations based on the tangent line. We remark that some well-known robustness measures are local, including Hampel, (1974)’s celebrated influence function (IF), and these local measures yield valuable insights for designing robust estimators and tests (Huber,, 1981). For more discussions behind the motivation for measuring local violations in robust statistics and how SLOPE yields valuable insights about robust study designs, see Sections 3 and 4.1. Second, from the definition of SLOPE, the unit of SLOPE inherits the unit of the target estimand, thereby respecting the investigator’s original choice of units for the target estimand. If the investigator wishes a unit-less SLOPE, a simple solution would be to transform the target estimand to be unit-less (e.g., in z-score units); see Section 3.1 for more discussions on interpreting SLOPE. Third, SLOPE depends on the parametrization of $Q_{O\mid X}^{\gamma}$ or equivalently, the sensitivity model for conditional exchangeability. Our sensitivity model has some benefits, but also carries some limitations (see Section 2.2, Section 4.1, and Remark 2). Regardless of the choice of the sensitivity model, we believe the high-level idea of SLOPE as a derivative-based summary of a sensitivity analysis can provide new and important insights about designing robust studies for generalization.
@@
Figure 1: Left: the data structure of our setup. Right: an illustration of SLOPE. The $y$-axis plots the target estimand/functional $\psi(\cdot)$ and the x-axis plots $\gamma$, which represents the magnitude of violating conditional exchangeability. The point $\gamma=0$ indicates no violation of conditional exchangeability.
@@ 1.2 Prior Works
SLOPE fits into the large literature on robust statistics. Specifically, SLOPE and Hampel, (1974)’s IF are related in that both SLOPE and IF use local derivatives to quantify robustness. But, Hampel, (1974)’s IF measures a local change in the target estimand $\psi$ due to contamination of a single data point whereas SLOPE measures a local change in $\psi$ due to violations of conditional exchangeability. Despite these differences, Section 4.1 reveals an interesting analytic and geometric connection between SLOPE and IF.
@@
Our work also fits into the literature on sensitivity analysis of conditional exchangeability for generalizability and transportability (Nguyen et al.,, 2017; Nie et al.,, 2021; Colnet et al.,, 2021; Dahabreh et al.,, 2022, 2023; Duong et al.,, 2023; Ek and Zachariah,, 2023; Huang,, 2024; Jin et al.,, 2024). What differentiates our paper from most existing works on sensitivity analysis for generalizability is that existing works studied sensitivity of estimators or tests for a specific value of the sensitivity parameter $\gamma\neq 0$. In contrast, our work studies sensitivity of study designs; as mentioned in our summary above, SLOPE is agnostic to how the target estimand $\psi$ is estimated and is a population-level characteristic.
@@
There have been some works on presenting simple, numerical summaries about the impact of violating key assumptions in different fields and we highlight some relevant examples. Gupta and Rothenhäusler, (2023) proposed the directional s-value, which quantifies the minimum amount of covariate shift that alters the sign of a target estimand, usually the mean. Rosenbaum, (2004) proposed design sensitivity, which is a scalar, odds-ratio based summary of the power of a test statistic in a sensitivity analysis. Andrews et al., (2017) considered local mis-specification of generalized method of moments (GMM) and their matrix-based measure, denoted as $\Lambda\in\mathbb{R}^{p\times p}$, measured the change in the estimator for GMM parameters. In missing data, Troxel et al., (2004) proposed a measure which quantifies the change in an estimator when there is a local violation of ignorable missingness. Ding and VanderWeele, (2016), Oster, (2019), Zhao, 2019b , and Cinelli and Hazlett, (2020) proposed scalar measures, which summarize the impact of unmeasured confounding in observational studies. Specifically, each proposed metrics that quantify the “minimum unmeasured confounding bias” necessary to alter the study’s conclusion under no unmeasured confounding. Except for Gupta and Rothenhäusler, (2023) and Rosenbaum, (2004) to some extent, all the above works measure robustness of estimators or tests rather than that of study designs. Also, except for Troxel et al., (2004), all these works do not use derivative-based measures of robustness.
Finally, we briefly mention another important line of work on robustness and sensitivity analysis when conditional exchangeability holds, but the distributions of shared characteristics $X$ of the two populations differ dramatically, especially in the observed data; this setting referred to as limited overlap (Stuart et al.,, 2011; Tipton,, 2014; Chen et al., 2023b, ; Huang,, 2025). Except for Huang, (2025), a key distinction between these works and our work is that limited overlap can be, in principle, checked from the observed data since $X$ is observed in both populations. In contrast, conditional exchangeability cannot be checked from the observed data since $O$ is not observed in the target population.
1.3 Organization of Paper and Notation
@@
The paper is organized as follows. Section 2 introduces the setup and the sensitivity model. Section 3 formally introduces SLOPE, its properties, and results on robust study designs. Section 4 discusses more insights and results, including the relationship between SLOPE and IF and estimating SLOPE. Section 5 showcases an application of SLOPE through a re-analysis of a multi-site experiment by Banerjee et al., (2015). Section 6 concludes with a discussion on practical considerations. Proofs and other results are in the Supplement.
@@
We define the notations that we use in the paper. For a population $P$ and random vectors $X_{1}$, $X_{2}$, we let $P_{X_{j}}$ be the marginal distribution of $X_{j}$ for $j=1,2$. Given $P_{X_{j}}$, we let $\textnormal{E}_{P_{X_{j}}}(\cdot)$ and $F_{P_{X_{j}}}$ be the expectation and cumulative distribution function (c.d.f.), respectively, under $P_{X_{j}}$. We let $P_{X_{1}\mid X_{2}}(\cdot\mid\cdot)$ be the conditional distribution of $X_{1}$ given $X_{2}$. Similarly, we let $\textnormal{E}_{X_{1}\mid X_{2}}(\cdot\mid\cdot)$, $F_{X_{1}\mid X_{2}}(\cdot\mid\cdot)$, $f_{X_{1}\mid X_{2}}(\cdot\mid\cdot)$ be the conditional distribution, conditional c.d.f., and conditional probability density function, respectively. We also let $\textnormal{E}_{X_{1}\mid X_{2}}(\cdot\mid x_{2})$ be the conditional expectation given a specific $X_{2}=x_{2}$. Throughout the paper, we assume sufficient regularity conditions for conditional distributions, conditional densities, and conditional expectations to exist; see Shao, (2008, Chapter 1) for the regularity conditions. Finally, for a probability distribution $Q$ defined on the same measurable space of $(X_{1},X_{2})$, we use $P_{X_{1}\mid X_{2}}\times Q_{X_{2}}$ to denote the joint distribution where $P_{X_{1}\mid X_{2}}\times Q_{X_{2}}(A\times B)=\int_{B}P_{X_{1}\mid X_{2}}(A\mid x_{2})dQ_{X_{2}}(x_{2})$ for $A\in\mathcal{B}$ and $B\in\mathcal{B}^{d}$ and $\mathcal{B}$ is a Borel $\sigma$-field with respect to the reals.
@@ 2 Setup
2.1 Goal in Generalization and Key Assumptions
Let $P_{O,X}$ and $Q_{O,X}$ be the joint distributions of random variables $(O,X)$ from a source population $P$ and a target population $Q$, respectively, where $O$ is a scalar and $X$ is a vector. The goal in generalization is to learn a functional (i.e., $\psi(\cdot)$) of the target distribution $Q_{O,X}$, which we denote as $\psi(Q_{O,X})$ and refer to as the target estimand or target functional, from (a) “full” data $(O,X)$ from the source distribution $P_{O,X}$ and (b) “partial” data $X$ from the target distribution $Q_{X}$; see Figure 1 for a visual illustration of the data setup.
Before we go any further, we make two brief remarks about the setup. First, while our exposition below considers a scalar functional $\psi$ (e.g., means, medians, average potential outcomes), all of our results extend to a low-dimensional, vector-valued $\psi$; see Section 3.1 and Section J.1 of the Supplement where we discuss the setting when $\psi$ is the regression coefficient of $O$ regressed on $X$ in the target population. Second, almost all results below are agnostic to how the data were sampled within each population, for instance by simple random sampling, i.i.d. sampling, or even adaptive sampling. Specifically, our results remain at the population level until Section 4.2, where we propose estimators of SLOPE.
Under the setup, the two most popular assumptions for identifying the target estimand $\psi$ (e.g., Cole and Stuart, (2010); Tipton, (2014); Kern et al., (2016); Dahabreh et al., (2019); Huang et al., (2023); Zeng et al., (2023); Degtiar and Rose, (2023)) are as follows.